# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v3

@@ 1 Introduction
[…]
@@ 2.1 Working Hypotheses
$H_{0}:{w}_{1:n}~{}\text{is written by a human}~{}~{}~{}~{}~{}~{}~{}H_{1}:{w}_{1:n}~{}\text{is written by a watermarked LLM}.$ (4)
@@ 2.2 Pivotal Statistics
$T_{h}(Y_{1:n}):=\left\{\begin{array}[]{ll}1&~{}\text{if }\sum_{t=1}^{n}h(Y_{t})\geq\gamma_{n,\alpha}\\ 0&~{}\text{if }\sum_{t=1}^{n}h(Y_{t})<\gamma_{n,\alpha}.\\ \end{array}\right.$ (10)
@@ 2.3 Class-Dependent Efficiency
Assume that at least $\gamma$-fraction of $\bm{P}_{1},\ldots,\bm{P}_{n}$ is in ${\mathcal{P}}_{1}$ with the rest being in ${\mathcal{P}}_{2}$, [^12] We do not need to know which are in the $\gamma$ fraction. where $10<\gamma<$ 1 and ${\mathcal{P}}_{1}\subset{\mathcal{P}}_{2}$. Then, for any $h$, we have
@@
When there is no prior at all about the LLMs, the associated class is the $(|{{\mathcal{W}}}|-1)$-dimensional simplex, which contains all possible NTP distributions and we denote by ${\mathrm{Simp}}({{\mathcal{W}}})$. This distribution class includes singular distributions—a distribution $\bm{P}$ such that $P_{{w}}=1$ for some token ${w}$—for which the pivotal statistic has the same distribution under the null and alternative in (9). Formally, this gives $R_{{\mathrm{Simp}}({{\mathcal{W}}})}(h)=0$, that is, the class-dependent efficiency of ${\mathrm{Simp}}({{\mathcal{W}}})$ is zero. Taking ${\mathcal{P}}_{1}={\mathcal{P}}\subset{\mathcal{P}}_{2}={\mathrm{Simp}}({{\mathcal{W}}})$, Proposition 2.1 shows $\lim\sup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}\leq\mathrm{e}^{-\gamma R_{{\mathcal{P}}}(h)}.$ This result implies that when only a fraction of NTP distributions is known to some extent, the Type II error is still governed by $R_{{\mathcal{P}}}(h)$ up to a multiplicative constant $\gamma$. The comparison between choices of the score functions $h$ can still be reduced to comparing the value of $R_{{\mathcal{P}}}(h)$.
@@
$-\inf_{\theta\geq 0}\log\biggl{[}\frac{1}{\theta}\biggl{\{}\frac{{\mathrm{e}}^{\theta(1-2\Delta)}+{\mathrm{e}}^{-\theta(1-2\Delta)}}{2}-{\mathrm{e}}^{-\theta}\biggr{\}}\biggr{]}.$
@@ 3 Application to the Gumbel-max Watermark
where =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution and $U\sim U(0,1)$.
@@ 3.1 Main Results
While it is now known that ${h_{\mathrm{gum},\Delta}^{\star}}$ is optimal in the sense of ${\mathcal{P}}_{\Delta}$-efficiency, it remains interesting to analyze the relative efficiency of other existing rules. In addition to the aforementioned $h_{\mathrm{ars}}$ and $h_{\mathrm{log}}$, we consider $h_{\mathrm{ind},\delta}(r)={\bf 1}_{\{r\geq\delta\}}$ for a threshold $10<\delta<$ 1, say, $\delta={\mathrm{e}}^{-1}$. [^14] According to the $\overline{R}_{\Delta}$-efficiency defined in (25), $\delta={\mathrm{e}}^{-1}$ achieves the largest $\overline{R}_{\Delta}$-efficiency when $\Delta\to 1$. See Lemma C.7 for details. This indicator function is included as it is perhaps the simplest score function detecting the distributional differences between $H_{0}$ and $H_{1}$.
@@
Due to the uniqueness of the Donsker–Varadhan representation, $\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{gum}})+\log\phi_{h}(\bm{P}_{\Delta}^{\star})$ is strictly larger than $-D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}_{\Delta}^{\star}})$ unless we take the log-likelihood ratio
@@ 4.1 Main Results
where $\Pi$ collects all permutations on ${{\mathcal{W}}}$, $a_{\pi,i}=\sum_{j=1}^{i}P_{t,\pi(j)}$, $B(v,r)=\{x\in[0,1]:|x-v|\leq r\}$ and ||$\cdot$|| ⋅ | represents the length of an interval.
@@
[…]
@@
Then, $\lim_{M\to\infty}\overline{R}_{\Delta}\left([{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}\right)=\infty.$
@@
Finally, the truncation in $[{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}$ is mainly for technical reasons, stemming from a proof artifact.
@@
For any function $h$ that is non-decreasing, Lipschitz-continuous on $[0,1]$, there is the lower bound:
@@
where the last identity holds because the two probability measures $\mu_{{\mathrm{dif}},0}$ and $\mu_{{\mathrm{dif}},\Delta}$ have different support, with $\mu_{{\mathrm{dif}},0}$ supported on $[0,1]$ and $\mu_{{\mathrm{dif}},\Delta}$ on $[0,1-\Delta]$ for $\Delta>0$. This supremum $\sup_{h}G(h)$ is obtained for $h$ as the log-likelihood ratio:
@@
[…]
@@
lim $\begin{split}&\limsup\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\left[\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{dif}})+\log\operatorname{\mathbb{E}}_{1,\bm{P}}[{\mathrm{e}}^{-h(Y^{\mathrm{dif}})}]\right]\\ =&\sup_{\Delta^{\prime}\geq\Delta}\limsup\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\left[\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{dif}})+\log\operatorname{\mathbb{E}}_{1,\bm{P}}[{\mathrm{e}}^{-h(Y^{\mathrm{dif}})}]\right].\end{split}$
[…]
@@ Appendix A Proof Supporting the General Framework
$T_{h}(Y_{1:n})=\left\{\begin{array}[]{ll}1&~{}\text{if}~{}\sum_{t=1}^{n}h(Y_{t})\geq\gamma_{n,\alpha}\\ 0&~{}\text{if}~{}\sum_{t=1}^{n}h(Y_{t})<\gamma_{n,\alpha},\\ \end{array}\right.$
@@
$\operatorname{\mathbb{E}}_{1,\bm{P}_{t}}\exp(-\theta h(Y_{t}))=\exp(\log\phi_{P_{t},h}(\theta))\leq\exp(\log\phi_{{\mathcal{P}},h}(\theta))$
@@
By definition, $\gamma_{n,\alpha}$ is the theoretical critical value that makes the detection rule $T_{h}$ have a given Type I error of $\alpha$, i.e.,
@@
By standard results from the large deviation theory [18, 65], if we set all the NTP distribution as $\bm{P}_{k}$, then
@@
$\leq\exp(-[\gamma R_{{\mathcal{P}}_{1},h}+(1-\gamma)R_{{\mathcal{P}}_{2},h}])$
@@ A.3 Difficulties of Finding the Optimal Pivot Statistic
To set up the stage, we introduce some notation. For a given pivotal statistic $Y$, we first denote $J(Y)=\sup_{h}R_{{\mathcal{P}}}(h)$ to be the ${\mathcal{P}}$ efficiency for that pivotal statistic. We deduce
@@
$=-\sup_{\bm{P}\in{\mathcal{P}}}\inf_{h}\left(\theta\operatorname{\mathbb{E}}_{0}h(Y)+\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-h(Y)}\right)$
@@
Here we recall the notation that ${w}$ refers to the token, and $\zeta$ refers to the pseudorandom numbers. We give examples to demonstrate that (i) the functional $J$ is not convex nor concave, and (ii) the constraint set $Y$ being a pivotal statistic is not a convex set.
@@
There is a singleton set $\{{\mathcal{P}}\}=\{\bm{P}\}$ with $Y_{1},Y_{2},Y_{3},Y_{4}$, each of which is a function of $({w},\zeta)$, such that
@@
Table 1 presents some important calculation results which complete the proof. ∎
@@
Table 1: CDFs under different hypotheses and the efficiency rates of the considered functions. $F_{i}$ represents the CDF under hypothesis $H_{i}$. The domain for all the CDFs is $r\in[0,1]$. Here ${w}=1$ if and only if $\frac{\log U_{1}}{P_{1}}\geq\frac{\log U_{2}}{P_{2}}$. Recall that we use $P_{1}=P_{2}=1/2$ in the proof of Proposition A.1.
@@
Define $Y_{1}=U_{{w}}$ and $Y_{2}=U_{1}$. It is clear that $Y_{1}$ and $Y_{2}$ are pivotal statistics, as their distributions are independent of the distribution of ${w}$ under the null hypothesis $H_{0}$. In the following, we aim to compute the null CDF for $\frac{Y_{1}+Y_{2}}{2}$. Under $H_{0}$, we have $w\sim\bm{P}=(P_{1},P_{2})$ and $w\perp(U_{1},U_{2})$. As a result, it follows that for any $r\in[0,1]$,
@@
We conclude that $\frac{Y_{1}+Y_{2}}{2}$ is not a pivotal statistic because its null CDF depends on the distribution of ${w}$, i.e., $\bm{P}=(P_{1},P_{2})$. ∎
@@ B.1 Analysis for the pivotal statistic $Y^{\mathrm{gum}}$
We assume $\zeta=(U_{w})_{{w}\in{{\mathcal{W}}}}$ with each $U_{{w}}$ i.i.d. from ${U}(0,1)$. We split the probability ${\mathbb{P}}(U_{{w}_{t}}\leq r)$ into the sum of probability of $|{{\mathcal{W}}}|$ disjoint events as follows:
@@
Suppose we have one generated token ($n=1$). Fix $t$ (and we drop the notation dependence on $t$). Consider the independence test regarding the token ${w}$ and the pseudorandom variable $\zeta$. Under $H_{0}$, ${w}$ and $\zeta$ are independent. Under $H_{1}$, ${w}$ and $\zeta$ are related by the decoder function ${\mathcal{S}}$. More precisely, we have the hypothesis testing problem:
@@
Assume $\nu:{{\mathbb{R}}}\to\overline{{{\mathbb{R}}}}$ is non-decreasing, right continuous, and satisfies ${\mathbb{P}}(\nu^{-1}(\log U_{1}/P_{1})\leq\nu^{-1}(\log U_{2}/P_{2}))=P_{1}/(P_{1}+P_{2})$ for $U_{1},U_{2}\overset{i.i.d.}{\sim}{U}(0,1)$ and any $P_{1},P_{2}>2$. Then $\nu^{-1}$ is strictly increasing, i.e., we have $\nu^{-1}(t_{1})<\nu^{-1}(t_{2})$ for any $t_{1}<t_{2}$.

### v4

@@ 1 Introduction
After generating text in the form of a token sequence, ${w}_{1}{w}_{2}\cdots{w}_{t-1}$, the (unwatermarked) LLM generates the next token ${w}_{t}$ according to a multinomial distribution $\bm{P}_{t}:=(P_{t,w})_{w\in{{\mathcal{W}}}}$ on the vocabulary $\mathcal{W}$}}, satisfying $\sum_{w\in{{\mathcal{W}}}}P_{t,w}=1$. [^2] As a convention, throughout this paper three subscripts appear: $\text{letter}_{t},\text{letter}_{{w}}$, and $\text{letter}_{t,{w}}$, which correspond to step $t$, token w{w}, and token w{w} at step $t$, respectively. We call $\bm{P}_{t}$ the next-token prediction (NTP) distribution at step $t$, and it depends on all prior generated tokens, the user-supplied prompt, as well as system prompts that are hidden from users [66, 52, 11].
@@ 2.1 Working Hypotheses
[…]
@@ 2.2 Pivotal Statistics
[…]
@@ 2.3 Class-Dependent Efficiency
Assume that at least $\gamma$-fraction of $\bm{P}_{1},\ldots,\bm{P}_{n}$ is in ${\mathcal{P}}_{1}$ with the rest being in ${\mathcal{P}}_{2}$, [^12] We do not need to know which are in the $\gamma$ fraction. where $0<\gamma<1$ and ${\mathcal{P}}_{1}\subset{\mathcal{P}}_{2}$. Then, for any $h$, we have
@@
When there is no prior at all about the LLMs, the associated class is the $(|{{\mathcal{W}}}|-1)$-dimensional simplex, which contains all possible NTP distributions and we denote by ${\mathrm{Simp}}({{\mathcal{W}}})$. This distribution class includes singular distributions—a distribution $\bm{P}$ such that $P_{{w}}=1$ for some token w{w}—for which the pivotal statistic has the same distribution under the null and alternative in (9). Formally, this gives $R_{{\mathrm{Simp}}({{\mathcal{W}}})}(h)=0$, that is, the class-dependent efficiency of ${\mathrm{Simp}}({{\mathcal{W}}})$ is zero. Taking ${\mathcal{P}}_{1}={\mathcal{P}}\subset{\mathcal{P}}_{2}={\mathrm{Simp}}({{\mathcal{W}}})$, Proposition 2.1 shows $\lim\sup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}\leq\mathrm{e}^{-\gamma R_{{\mathcal{P}}}(h)}.$ This result implies that when only a fraction of NTP distributions is known to some extent, the Type II error is still governed by $R_{{\mathcal{P}}}(h)$ up to a multiplicative constant $\gamma$. The comparison between choices of the score functions $h$ can still be reduced to comparing the value of $R_{{\mathcal{P}}}(h)$.
@@
$\inf_{\theta\geq 0}\log\biggl{[}\frac{1}{\theta}\biggl{\{}\frac{{\mathrm{e}}^{\theta(1-2\Delta)}+{\mathrm{e}}^{-\theta(1-2\Delta)}}{2}-{\mathrm{e}}^{-\theta}\biggr{\}}\biggr{]}.$
@@ 3 Application to the Gumbel-max Watermark
where $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution and $U\sim U(0,1)$.
@@ 3.1 Main Results
While it is now known that $h_{\mathrm{gum},\Delta}^{\star}$} is optimal in the sense of ${\mathcal{P}}_{\Delta}$-efficiency, it remains interesting to analyze the relative efficiency of other existing rules. In addition to the aforementioned $h_{\mathrm{ars}}$ and $h_{\mathrm{log}}$, we consider $h_{\mathrm{ind},\delta}(r)={\bf 1}_{\{r\geq\delta\}}$ for a threshold $0<\delta<1$, say, $\delta={\mathrm{e}}^{-1}$. [^14] According to the $\overline{R}_{\Delta}$-efficiency defined in (25), $\delta={\mathrm{e}}^{-1}$ achieves the largest $\overline{R}_{\Delta}$-efficiency when $\Delta\to 1$. See Lemma C.7 for details. This indicator function is included as it is perhaps the simplest score function detecting the distributional differences between $H_{0}$ and $H_{1}$.
@@
Due to the uniqueness of the Donsker–Varadhan representation, $\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{gum}})+\log\phi_{h}(\bm{P}_{\Delta}^{\star})$ is strictly larger than $D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}_{\Delta}^{\star}})$ unless we take the log-likelihood ratio
@@ 4.1 Main Results
where $\Pi$ collects all permutations on $\mathcal{W}$}}, $a_{\pi,i}=\sum_{j=1}^{i}P_{t,\pi(j)}$, $B(v,r)=\{x\in[0,1]:|x-v|\leq r\}$ and $|\cdot|$ represents the length of an interval.
@@
where $\liminf$ denotes the large vocabulary size limit where $|{{\mathcal{W}}}|\to\infty$ and $\log|{{\mathcal{W}}}|\cdot{\varepsilon}_{|{{\mathcal{W}}}|}\to 0$. In plain words, $\overline{R}_{\Delta}(h)$ quantifies the efficiency of any score function $h$ in the limit of large vocabulary size $|{{\mathcal{W}}}|\to\infty$, assuming that the distribution of the underlying LLM predominantly focuses on a single token.
@@
Then, $\lim_{M\to\infty}\overline{R}_{\Delta}\left([{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}\right)=\infty$ where $x]_{[-M,M]}:=\min\{\max\{x,-M\},M\}$ denotes the clipping operator that restricts $x$ to the interval [M,M][-M,M].
@@
Finally, the truncation in ${h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}$ is mainly for technical reasons, stemming from a proof artifact.
@@
For any function $h$ that is non-increasing, Lipschitz-continuous on $[0,1]$, there is the lower bound:
@@
where the last identity holds because the two probability measures $\mu_{{\mathrm{dif}},0}$ and $\mu_{{\mathrm{dif}},\Delta}$ have different support, with $\mu_{{\mathrm{dif}},0}$ supported on $[0,1]$ and $\mu_{{\mathrm{dif}},\Delta}$ on $0,1-\Delta]$ for $\Delta>0$. This supremum $\sup_{h}G(h)$ is obtained for $h$ as the log-likelihood ratio:
@@
We wish to swap the order of $\sup_{\Delta^{\prime}\geq\Delta}$ and $\limsup$ in (28) to reach the conclusion of Lemma 4.3. Lemma 4.4 justifies such a swap as valid.
@@
$\begin{split}&\limsup\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\left[\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{dif}})+\log\operatorname{\mathbb{E}}_{1,\bm{P}}[{\mathrm{e}}^{-h(Y^{\mathrm{dif}})}]\right]\\ =&\sup_{\Delta^{\prime}\geq\Delta}\limsup\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\left[\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{dif}})+\log\operatorname{\mathbb{E}}_{1,\bm{P}}[{\mathrm{e}}^{-h(Y^{\mathrm{dif}})}]\right].\end{split}$
By swapping $\sup_{\Delta^{\prime}\geq\Delta}$ and $\limsup$ in (28), we obtain
@@ Acknowledgments
Supplementary Material
This Supplementary Material contains the remaining proofs and technical details. The proof that supports the general framework is collected in Section A. The proofs about the Gumbel-max watermark are presented in Section B. Section C includes the proofs of results for the inverse transform watermark. Section D contains experiment details.
@@ Appendix A Proof Supporting the General Framework
[…]
@@
$\operatorname{\mathbb{E}}_{1,\bm{P}_{t}}\exp(-\theta h(Y_{t}))=\exp(\log\phi_{P_{t},h}(\theta))\leq\exp(\log\phi_{{\mathcal{P}},h}(\theta)).$
@@
By definition, $\gamma_{n,\alpha}$ is the theoretical critical value that makes the detection rule $T_{h}$ have a given Type I error of $\alpha$, that is,
@@
By standard i.i.d. results from the large deviation theory [18, 65], if we set all the NTP distribution as $\bm{P}_{k}$, then
@@
$\leq\exp(-[\gamma R_{{\mathcal{P}}_{1},h}+(1-\gamma)R_{{\mathcal{P}}_{2},h}]).$
@@ A.3 Difficulties of Finding the Optimal Pivot Statistic
To set up the stage, we introduce some notation. For a given pivotal statistic $Y$, we first denote $J(Y)=\sup_{h}R_{{\mathcal{P}}}(h)$ to be the $\mathcal{P}$}-efficiency for that pivotal statistic. We consider the simplest case where $\mathcal{P}$} contains only a single element (so that we can exchange the order of $\sup_{\bm{P}\in{\mathcal{P}}}$ and $\inf_{h}$). We deduce that
@@
$=-\sup_{\bm{P}\in{\mathcal{P}}}\inf_{h}\left(\operatorname{\mathbb{E}}_{0}h(Y)+\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-h(Y)}\right)$
@@
Here, we recall the notation that w{w} refers to the token, and $\zeta$ refers to the pseudorandom numbers. We give examples to demonstrate that (i) the functional $J$ is neither convex nor concave, and (ii) the constraint set consisting of all $Y$ that form a pivotal statistic is not convex.
@@
There is a singleton set ${\mathcal{P}}=\{\bm{P}\}$ with $Y_{1},Y_{2},Y_{3},Y_{4}$, each of which is a function of $({w},\zeta)$, such that
@@
Table 1 presents some important calculation results that complete the proof. ∎
@@
Table 1: CDFs under different hypotheses and the efficiency rates of the considered functions. $F_{i}$ represents the CDF under hypothesis $H_{i}$. The domain for all the CDFs is $r\in[0,1]$. Here w=1{w}=1 if and only if $\frac{\log U_{1}}{P_{1}}\geq\frac{\log U_{2}}{P_{2}}$. Recall that we use $P_{1}=P_{2}=1/2$ in the proof of Proposition A.1.
@@
Define $Y_{1}=U_{{w}}$ and $Y_{2}=U_{1}$. It is clear that $Y_{1}$ and $Y_{2}$ are pivotal statistics, as their distributions are independent of the distribution of w{w} under the null hypothesis $H_{0}$. In the following, we aim to compute the null CDF for $\frac{Y_{1}+Y_{2}}{2}$. Under $H_{0}$, we have $w\sim\bm{P}=(P_{1},P_{2})$ and $w\perp(U_{1},U_{2})$. As a result, it follows that for any $r\in[0,1]$,
@@
We conclude that $\frac{Y_{1}+Y_{2}}{2}$ is not a pivotal statistic because its null CDF depends on the distribution of w{w}, that is, $\bm{P}=(P_{1},P_{2})$. ∎
@@ B.1 Analysis for the pivotal statistic $Y^{\mathrm{gum}}$
We assume $\zeta=(U_{w})_{{w}\in{{\mathcal{W}}}}$ with each $U_{{w}}$ i.i.d. from U(0,1){U}(0,1). We split the probability ${\mathbb{P}}(U_{{w}_{t}}\leq r)$ into the sum of probability of $|{{\mathcal{W}}}|$ disjoint events as follows:
@@
Suppose we have one generated token ($n=1$). Fix $t$ (and we drop the notation dependence on $t$). Consider the independence test regarding the token w{w} and the pseudorandom variable $\zeta$. Under $H_{0}$, w{w} and $\zeta$ are independent. Under $H_{1}$, w{w} and $\zeta$ are related by the decoder function $\mathcal{S}$}. More precisely, we have the hypothesis testing problem:
@@
Assume $\nu:{{\mathbb{R}}}\to\overline{{{\mathbb{R}}}}$ is non-decreasing, right continuous, and satisfies ${\mathbb{P}}(\nu^{-1}(\log U_{1}/P_{1})\leq\nu^{-1}(\log U_{2}/P_{2}))=P_{1}/(P_{1}+P_{2})$ for $U_{1},U_{2}\overset{i.i.d.}{\sim}{U}(0,1)$ and any $P_{1},P_{2}>2$. Then $\nu^{-1}$ is strictly increasing, that is, we have $\nu^{-1}(t_{1})<\nu^{-1}(t_{2})$ for any $t_{1}<t_{2}$.